Calculate E° for the following voltaic cells: a) MnO2(s) + 4H^+(aq) + 2l^- = Mn^2(aq) + 2H2O + l2(s) b) 2S(s) + 4OH(aq)=O2(g) + 2S^2-(aq) + 2H2O
1
Expert's answer
2020-12-17T07:34:19-0500

"a) MnO_{2(s)} + 4H^+_{(aq)} + 2I^- = Mn^{2+}_{(aq)} + 2H_2O + I_{2(s)}"


Reduction

MnO2 --> Mn2+ (E = 1.23V)

Oxidation

2I- --> I2 (E = -0.54V)


E° = E(ox) + E(red) = -0.54 +1.23 = 0.69V


"b) 2S_{(s)} + 4OH^-_{(aq)} \to O_{2(g)} + 2S^{2-}(aq) + 2H_2O"


Reduction

2S --> 2S2- (E = 0.14V)

Oxidation

4OH- --> O2 + 2H2O (E = -0.401V)


E° = E(ox) + E(red) = -0.401 +0.14 = -0.261V